State of Economy Information V Probability of State of Economy Returns if State Occurs Stock II Boom Normal Wow 0.06 0.69 0.2Use INFORMATION V on stocks I and II: The market risk premium is 8 percent, and the risk-free rate is 3.6 percent. The beta of stock I is BLANK and the beta of stock II is BLANK.

  • A. 2.08; 2.47
  • B. 2.08; 2.76
  • C. 3.21; 3.84
  • D. 4.47; 3.89
  • E. 4.03; 3.71
Answer #1

Ans:- In this question, we need to find the beta of Stock I and II by Using Information V

The expected return is given by W1*R1+W2*R2+............WN*RN, where W is the weights or Probability, and R is the rate of return.

Expected return of Stock I = 0.06*0.15%+0.69*35%+0.25*43% = 35.80%

Expected return of Stock II =0.06*-35%+0.69*35%+0.25*45% = 33.30%

Now we will use the CAPM equation to find the Beta

By CAPM, Expected return = Risk-free rate + Beta * Market Risk Premium

or Beta = ( Expected return - Risk-free rate ) / Market Risk Premium

Risk-free rate = 3.6% and Market risk premium = 8%

Beta of stock I = ( 35.80% - 3.6% ) / 8% = 4.03 (approx)

Beta of Stock II = ( 33.30% - 3.6% ) / 8% = 3.71 (approx).

Therefore, Beta of Stock I is 4.03, and Beta of Stock II is 3.71. option E is the right answer.

Note:- If this answer helps you pls give thumbs up.